AziTech is considering the design of a new CPU for its new model of computer systems for
2021. It is considering choosing between two (2) CPU (CPUA and CPUB) implementations
based on their performance. Both CPU are expected to have the same instruction set
architecture. CPUA has a clock cycle time of 60 ns and CPUB has a clock cycle time of 75
ns. The same number of a particular instruction type is expected to be executed on both CPUs
in order to determine which CPU would executes more instructions. CPUA is able to execute
2MB of instructions in 5*106 clock cycles. CPUB executes the same number of instructions in
3*106 clock cycles.
1
Expert's answer
2020-06-26T04:26:41-0400

MIPs (million instructions per second) is the general measurement or benchmark of how many instructions a processor can handle in a single second. 

1.

CPUA is better as it executes 2MB of instructions in 5*106 clock cycles as the MIPS looks the execution of number of instructions per second. Whereas CPUB executes same number of instructions in 3*106 clock cycles clearly shows that CPUA is better on the basis of MIPS performance metrics.

2. 

CPU time (or CPU Execution time) is the time between the start 

and the end of execution of a given program. This time 

accounts for the time CPU is computing the given program, 

including operating system routines executed on the program’s 

behalf, and it does not include the time waiting for I/O and 

running other programs.

CPUA executed time is 1/cpu_time

1/5*106 = 0.00188679245 seconds

CPUB execution time is

1/cpu_time

1/5*106 = 0.00188679245 seconds

CPUB execution time is 1/cpu_time

1/3*106 = 0.00314465409 seconds


CPUB is faster than CPUA according to their execution time.